OpenStudy (anonymous):

Can help with 41^2 + 28^2 = c^2

OpenStudy (anonymous):

c=49.65

OpenStudy (anonymous):

41^2= 1681 28^2=784 1681+784=2465 Square root of 2465= 49.649

OpenStudy (anonymous):

thanks